OpenStudy (thephysicsman):

A box contains 10 balls numbered from 1 to 10 inclusive. If Ann removes a ball at random and replaces it, and then Jane removes a ball at random, what is the probability that both women removed the same ball?

OpenStudy (thephysicsman):

@DanJS

OpenStudy (3mar):

0.01

OpenStudy (thephysicsman):

hey @3mar can you explain that?

OpenStudy (sshayer):

\[\frac{ 1 }{ 10 }\times \frac{ 1 }{ 10 }=?\]

OpenStudy (thephysicsman):

\[\frac{ 1 }{ 100 }\]

OpenStudy (thephysicsman):

it's saying the answer is 1/10

OpenStudy (3mar):

Sorry it will be 0.1 Because the probability of the first one is 0.1 and the second will be the same 0.1 as it is the same ball.

OpenStudy (3mar):

If you got it, tell me.

OpenStudy (thephysicsman):

alright so if one person takes out a ball and puts it back it's 1/10

OpenStudy (thephysicsman):

then there are still 10 balls left. I get that part

OpenStudy (3mar):

Good.

OpenStudy (thephysicsman):

if the other person takes it out it's also 1/10. but I don't understand is that the probability of selecting the same ball wouldn't that mean 1/10*1/10

OpenStudy (3mar):

The second person will choose one ball of the 10 balls, the probability will be 0.1

OpenStudy (sshayer):

for each it is 1/10

OpenStudy (3mar):

sshayer is right.

OpenStudy (thephysicsman):

maybe i'm over thinking this. @3mar I guess i'm thinking about it as in the probability of selecting the same ball but we wouldn't multiply the probabilities because the women are putting the same ball back.

OpenStudy (3mar):

Yes It is 0.1 for each one, not meaning that 0.1*0.1